What of the following statements is/are correct?

  1. The current in one wire create electric field, which excerts a force on the second wire with the current;
  2. The current in one wire create magnetic field, which excerts a force on the second wire with the current;
  3. The force between the wires is inverse-proportional to the distance between wires;
  4. The force between the wires is inverse-proportional to the squared distance between wires;
  5. A and D;
  6. B and C.

Final answer too, please
Answer #1

Magnetic force between two current carrying wire is given by:

F/L = u0*i1*i2/(2*pi*d)

And force will be attractive if both wires have current in same direction and Force will be repulsive if both wires have

current in opposite direction

Given that: F = Force of attraction on each wire = 1.5 mN = 1.5*10^-3 N

i1 = 25 A, i2 = 15 A

L = length of wire = 1.0 m

So,

d = distance between both wires = u0*i1*i2*L/(2*pi*F)

d = 4*pi*10^-7*25*15*1.0/(2*pi*1.5*10^-3)

d = 0.05 m = 5.0 cm

Since both wires are attracting each other, So current in them is in same direction.

Part 2.

We know that Force on a current carrying wire is given by:

F = i1*L*B

And magnetic field generated by another wire is given by:

B = u0*i2/(2*pi*d), So

F = u0*i1*i2*L/(2*pi*d)

So The current in one wire create magnetic field, which excerts a force on the second wire with the current.

Also we can see from above that The force between the wires is inverse-proportional to the distance between wires.

So Correct option is F. (B and C)

Let me know if you've any query.